169cm^3 of hydrogen were collected over water at 20°c and 755 mmHg. The saturated vapour pressure of water at 20°c is 17 .5 mmHg. What will be the volume of the hydrogen at s.t.p


1
Expert's answer
2022-03-14T18:18:06-0400

Volume of hydrogen = 169 cm3 =V1, Temperature of hydrogen gas collected = 20oC= T1, Pressure of gas = 755 = P1

At stp , Temperature = 25oC=T2 and Pressure = P2= 760 mm Hg

Using ideal gas equation


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


"\frac{755×169}{293}=\frac{760×V_2}{298}"


"V_2= 170.7 cm^3"


Volume of gas at STP is =170.7 cm3